Opened 10 years ago

Closed 9 years ago

Last modified 3 years ago

#269 closed flight processing (fixed)

GB08/15, flight day 278a, Eaves Wood

Reported by: mark1 Owned by: knpa
Priority: alpha 4 low Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by emca)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB08_15-2009_278a_Eaves_wood

Data arrived from ARSF via SATA disk L on 7th October 2009.

Scientific objective:/users/rsg/arsf/arsf_data/2008/ARSF_Applications/GB_and_NL/GB08-15 Science Case.pdf

Priority: alpha-4-Low

PI: A. Blackburn

Sensors:

  • Eagle (02/06/2010)
  • Hawk (02/06/2010)
  • Leica LIDAR (02/06/2010)

Change History (20)

comment:1 Changed 10 years ago by mark1

Still awaiting flight logsheet.

DEM creation threw errors but created a DEM. Check it is good when processing.

comment:2 Changed 10 years ago by mark1

logsheet arrived and unpacked

comment:3 Changed 10 years ago by mark1

Have renamed hawk files from 278* and Hawk278* to SWIR278*

comment:4 Changed 10 years ago by mggr

Re: #281 (Hawk frame rates are double what they should be), the following lines are affected and the .hdr file will need to be edited to halve the fps number:

uk/GB08_15-2009_278a_Eaves_wood/hawk/SWIR278-09-10.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_15-2009_278a_Eaves_wood/hawk/SWIR278-09-11.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_15-2009_278a_Eaves_wood/hawk/SWIR278-09-1.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_15-2009_278a_Eaves_wood/hawk/SWIR278-09-2.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_15-2009_278a_Eaves_wood/hawk/SWIR278-09-3.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_15-2009_278a_Eaves_wood/hawk/SWIR278-09-4.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_15-2009_278a_Eaves_wood/hawk/SWIR278-09-5.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_15-2009_278a_Eaves_wood/hawk/SWIR278-09-6.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_15-2009_278a_Eaves_wood/hawk/SWIR278-09-7.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_15-2009_278a_Eaves_wood/hawk/SWIR278-09-8.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_15-2009_278a_Eaves_wood/hawk/SWIR278-09-9.hdr: predicted length half recorded length, frame rate doubling problem?

comment:5 Changed 10 years ago by emca

  • Owner set to emca
  • Status changed from new to assigned

comment:6 Changed 10 years ago by emca

Renamed the eagle files from 278* to VNIR278*

Halved the fps number on the hawk header files. No sync for any of the hawk files.

comment:7 Changed 10 years ago by emca

LIDAR processing complete and ready for delivery check at
~airborne/workspace/GB08_15-2009_278a_Eaves_wood/delivery/20100525/GB08_15

comment:8 Changed 10 years ago by emca

Sct offsets:

Line No Eagle Hawk
1 +0.02 -5.15
2 -0.04 -4.67
3 +0.04 -5.68
4 +0.02 -4.96
5 +0.01 -4.35
6 +0.01 -4.69
7 -0.04 -4.55
8 -0.02 -5.46
9 +0.02 -4.92
10 +0.03 -4.75
11 -0.01 -5.61

comment:9 Changed 10 years ago by emca

No dark frames from Hawk data - processed using dark frames from GB09/05 278c line 09

Hyperspectral processing complete and ready for delivery check at
~airborne/workspace/GB08_15-2009_278a_Eaves_wood/delivery/20100527/GB08_15

comment:10 Changed 10 years ago by mark1

LIDAR delivery checked. All looks good. Ready to deliver.

comment:11 Changed 10 years ago by mark1

Hyperspectral delivery checked. All data looks good, ready to deliver.

comment:12 Changed 10 years ago by emca

  • Description modified (diff)

Hyperspectral and LiDAR data sent to PI 02/06/2010

comment:13 Changed 9 years ago by emca

  • Component changed from Processing: general to Archiving

All data has been processed and rsync'd to
~arsf/arsf_data/2009/flight_data/uk/GB08_15-2009_278a_Eaves_wood

Ready for archiving.

comment:14 Changed 9 years ago by knpa

  • Owner changed from emca to knpa

Beginning archiving.

comment:15 Changed 9 years ago by knpa

No webcam data for this project.

comment:16 Changed 9 years ago by knpa

Archiving preparation completed, awaiting NEODC transfer.

comment:17 Changed 9 years ago by knpa

Downloaded and backed up by NEODC.

Moved to archive disk.

comment:18 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version.

Closing ticket.

comment:19 Changed 3 years ago by lah

Archiving

Tidied and ready for upload to NEODC.

comment:20 Changed 3 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.